# The behavioral spillover effect: Modeling behavioral interdependencies in multi-pathogen dynamics

### v1

@@ 1 Introduction
Most studies focus on modeling the spread of a single virus, such as influenza or coronavirus [10, 11, 12]; however, multiple pathogens often co-circulate, with their relative dominance shifting over time [13]. For example, influenza and respiratory syncytial virus frequently overlap during seasonal epidemics, with higher prevalence observed in colder months [14]. In contrast, Lyme disease is more common during warmer seasons when outdoor activities increase [15], and rhinoviruses peak in the spring and fall [13]. This raises an important question: to what extent are the dynamics of multiple co-circulating pathogens independent of each other? This question became especially important during early months of the COVID-19 pandemic when a large fraction of public health capacities was devoted to treating COVID-19 patients [16].
@@ 1.1 Cross immunity
The phenomenon of cross-immunity and evolution of various strains and mutations have been explored through mathematical modeling, e.g. [22, 23, 27, 28]. For example, Bhattacharyya et al. [22] examine cross-immunity among respiratory viruses, focusing specifically on RSV, HPIV, and hMPV. Using compartmental models, they analyze patterns of laboratory-confirmed infections from the western United States between 2002 and 2014, arguing that closely related paramyxoviruses can suppress each other’s transmission through immune responses triggered by prior infections. In foundational work, Gog and Grenfell [23], showed how cross-immunity among multiple pathogens can produce a range of complex, yet realistic behavior.
@@
1.2 Human behavior
Recent advancements in epidemic modeling have led to various methods of formulating human response in epidemic models. LeJeune et al. [9] offer a review of such approaches, primarily distinguishing between exogenous and endogenous approaches to human behavior formulation. While the former includes external changes of parameter values, the latter formulates contact rate, infectivity, and shifts across compartments as a function of the state of the disease. For example, LeJeune et al. [35] and Rahmandad et al. [8] construct and analyze the endogenous SEIRb model (with b standing for behavior), where the parameter for infectivity is a function of disease prevalence. Infectivity declines as the number of cases increases (with a delay), representing delayed adjustment in human risk perception and behavior. LeJeune et al. (2024) [35] showed how such formulations allow greater versatility within models for capturing realistic dynamics of disease waves, and tested the model against data from 50 US states and Washington, D.C. Additional endogenous approaches include representing the susceptible population in different compartments depending on their compliance with health measures in response to the state of the disease [36, 37] as well as modeling human behavior as a payoff maximizer by formulating human response as a function of economic costs associated with decreasing activities versus the likelihood of infection or death [38, 39]. Other complementary approaches include explicit formulation of the spread of fear, change in social norms, and adherence fatigue regarding NPI compliance along with change in the state of the disease [40, 41, 42].
While these works highlight recent advancements in the consideration of cross-immunity and human behavior in epidemiological models, we highlight here a critical gap which persists in the literature. On one hand, models of multi-pathogen dynamics often lack behavioral mechanisms; on the other hand, most behavioral epidemic models focus on single-pathogen contexts, overlooking how individual and collective responses influence outbreak dynamics. Particularly, compliance with NPIs can simultaneously influence the spread of multiple pathogens with similar transmission mechanisms.
@@ 1.3 Appearance in data
In real-world settings, public behavior is rarely pathogen-specific. For instance, increased mask-wearing or social distancing in response to one disease can inadvertently suppress the transmission of others. This was clearly observed during the COVID-19 pandemic, when widespread NPI substantially reduced flu cases [17, 18, 19]. As shown in Figure 1, the United States has historically experienced annual flu seasons, typically peaking in late fall or early winter, until the onset of the recent COVID-19 pandemic. During the surge in COVID-19 cases in 2020 and 2021, influenza cases dramatically declined, with very few cases of Influenza A or B. Interestingly, as COVID-19 was brought under control and transitioned into an endemic phase, seasonal flu patterns re-emerged. In fact, Flu A cases surged during the 2022–2023 season. This pattern is often attributed to behavioral factors: compliance with NPIs not only curbed COVID-19 transmission but also suppressed flu outbreaks. Once such measures were relaxed, flu incidence rebounded.
@@ 1.4 Beyond specific diseases
Similar dynamics are likely to arise in a variety of contexts and among pathogens with comparable transmission characteristics. For example, when individuals increase hand washing in response to a surge in one particular disease, they may also inadvertently reduce the transmission of other pathogens that spread through similar pathways. As such, modeling frameworks that incorporate behavioral coupling between co-circulating pathogens can offer unique insights into epidemic dynamics. These models can help answer the following important questions: How do general patterns of disease spread emerge when multiple pathogens are active and NPIs targeting one also reduce the transmission of another? How might these patterns change if such spillover effects are imperfect or only partially suppress the spread of a second disease? Can we identify ranges of reproductive numbers where one pathogen dominates or where multiple pathogens can coexist? Addressing these questions is essential for improving our understanding of—and response to—the complex interplay between human behavior and the concurrent spread of multiple infectious diseases.
In this paper, we study the short-term and long-term dynamics of two pathogens spreading simultaneously, where NPIs targeting one pathogen can indirectly influence the spread of others—a phenomenon we term behavioral spillover. We present a mathematical model of two interacting pathogens and analyze their outbreak dynamics. The interaction is solely via human behavior; that is, no cross-immunity is considered, but we model how human risk response to one can fully or partially affect the spread of the other pathogen. We postulate that, in the short run, there are two regimes – persistence by only one pathogen or coexistence of both – and examine tipping points for change in patterns. Such equilibrium is influenced by the basic reproduction numbers of the pathogens. In the long run, we observe a likelihood of periodic shift in dominance between pathogens.
@@ 2 Materials and Methods
The core of our model structure is the simple SIR model with two modifications that include behavioral response and waning immunity. We refer to this family of models as SIRSb models to indicate the return of those in the recovered compartment to the susceptible compartment, and the inclusion of behavioral feedback affecting infectivity [35].
@@
where $S_{i}$, $I_{i}$, and $R_{i}$ represent the proportion of individuals in the total population who are susceptible, infectious, and recovered, respectively, and $\beta_{i}(\cdot)$ is the transmission rate, which may depend on other variables. Here, $\widetilde{I}_{i}$ represents perceived infectious individuals after average lag of $\tau_{P}$. Susceptible individuals become infected with disease $i$ through interactions with infectious individuals of disease $i$ and recover after an average period of $\tau_{I}$ days. Once recovered, individuals lose immunity and are once again susceptible after on average $\tau_{R}$ days. Parameter values used for the two diseases ($A$ and $B$) are given in Table 1.
@@
which represents the effect of response (e.g., decline in contact rate) on transmission rate of $\beta_{i}$. Here, $m_{i}$ is formulated such that increases in prevalence ($I$), tracked through a lagged compartment of the infectious population, $\widetilde{I}_{i}$, lead to a decline in $m_{i}$. As prevalence $I_{i}$ increases, perceived risk of $\widetilde{I}_{i}$ (modeled by an exponential delay of prevalence) also increases. Higher perceived risk leads to stronger reaction, decreasing $e_{i}$, and consequently decreasing $\beta_{i}$.
Symbol Description Value
$\mathcal{R}_{0,A}$ basic reproduction number of disease AA 3
$\mathcal{R}_{0,B}$ basic reproduction number of disease $B$ [1,3]
$\beta_{0,A}$ infectivity rate of disease AA $\frac{\mathcal{R}_{0,A}}{\tau_{I}}$
$\beta_{0,B}$ infectivity rate of disease $B$ $\frac{\mathcal{R}_{0,B}}{\tau_{I}}$
$\tau_{I}$ infection period 7 days
$\tau_{p}$ delay to adjust risk perception 30 days
$\tau_{R}$ immunity period 100 days
$k$ sensitivity to risk 100
$s$ spillover constant [0,1]
$S_{A}(0)$ initial proportion of disease-$A$ susceptible individuals 0.9999
$S_{B}(0)$ initial proportion of disease-$B$ susceptible individuals 0.9999
$I_{A}(0)$ initial proportion of disease-$A$ infectious individuals 0.0001
$I_{B}(0)$ initial proportion of disease-$B$ infectious individuals 0.0001
$R_{A}(0)$ initial proportion of disease-$A$ recovered individuals 0
$R_{B}(0)$ initial proportion of disease-$B$ recovered individuals 0
$\widetilde{I}_{A}(0)$ initial proportion of disease-$A$ perceived infectious individuals 0
$\widetilde{I}_{B}(0)$ initial proportion of disease-$B$ perceived infectious individuals 0
@@ 2.1 Transmission rate
We consider three scenarios, differing with respect to the level of behavioral spillover. Mathematically, this leads to differences in the multipliers $m_{A}$ and $m_{B}$ in the formulation of the transmission rate ($\beta_{i}(\cdot)$) in the SIRS model given in System (LABEL:eqn:SIRS). The specific system of equations for each scenario, with the relevant formulation of transmission, are included in Appendix A1.1 for the reader’s reference.
@@ 2.1.1 Two independent diseases
For the case of two pathogens which are not behaviorally coupled and with no cross-immunity, we assume that the transmission rate of each pathogen is formulated independently as
@@
where $\beta_{0,i}$ is the infectivity of disease $i$ and $m_{i}$ the multiplier of disease $i$.
@@ 2.1.2 Two behaviorally-coupled diseases
As a first scenario, we assume the simple case of perfect spillover. This represents a situation where a proportional decline in transmission of disease $A$, due to behavioral response, results in the same proportional decline in the transmission of disease $B$, and vice versa. This happens when the two diseases transmit in an identical manner. Behavioral response to one can alter the spread of the other, such that transmission of each disease is affected by both $m_{A}$ and $m_{B}$. Thus we modify the transmission rate of disease $i$, $\beta_{i}$, to be
@@
It is important to note that this formulation encapsulates both no spillover (independent diseases) and perfect spillover. For perfect spillover ($s=1$), the overall effect of response to $B$ on the spread of disease $A$ is $m_{B}$, and for no spillover ($s=0$), the overall effect of response to disease $B$ on the spread of $A$ is constant at one. Table 2 summarizes all the formulations for $\beta_{i}(\cdot)$ with $i\in\{A,B\}$. We note that in model formulation and analytical results, we present no spillover ($s=0$) and perfect spillover ($s=1$) before discussing imperfect spillover ($0<s<1$), while for numerical results, we present results ordered according to increasing spillover.
@@ 3 Results
We begin by numerically simulating the dynamics of our system with a range of spillover – none, perfect and imperfect (Section 3.1). Following our numerical examination, we determine the existence and stability of equilibria of the model under each spillover scenario (Section 3.2) and assess the structural and practical identifiability of the model (Section 3.3).
@@ 3.1 Numerical results
To examine the dynamics of our behaviorally-coupled system, we numerically simulate our system in MATLAB [45] using ode45 and observe the dynamics over short (one year) and long (multiple years) time scales. Figure 3 shows results from nine simulation experiments which vary the level of spillover and the basic reproduction number of disease $B$ (with the basic reproduction number of disease $A$ fixed at 3). A helpful quantity for explaining the behavior in the system is the effective reproduction number, $\mathcal{R}_{e}$, which tells if the disease is expanding or contracting at the current time. For our two disease system, we define the effective reproduction number of disease $i$, $\mathcal{R}_{e,i}$, as
@@
Without spillover (under our assumption $\mathcal{R}_{0,A}>\mathcal{R}_{0,B}>1$), there are two regimes of behavior: (i) diseases $A$ and $B$ co-exist but disease $A$ prevalence always remains above disease $B$ prevalence and (ii) diseases $A$ and $B$ co-exist but the diseases alternate whose prevalence is higher. As depicted in the left column of Figure 3 (Panels adg), both diseases persist as we assume the basic reproduction number of each disease independently is above one. The endemic level of the disease is dictated by its basic reproduction number. Furthermore, the dynamics are qualitatively similar for the two diseases and are quantitatively identical when the diseases have the same traits. As cases grow for disease $i$, individuals respond, leading to reductions in the transmission rate and a drop in $\mathcal{R}_{e,i}$. Ultimately, $\mathcal{R}_{e,i}$ settles to one such that cases reach a non-zero equilibrium (under our assumption $\mathcal{R}_{0,i}>1$). As the basic reproduction number of disease $B$ is lowered, the initial peak occurs later and reaches a lower level (Figure 3adg, red lines as $\mathcal{R}_{0,B}$ is lowered from 2.9 to 2 to 1.3). This is because the natural disease infectivity is lower, reducing the transmission rate even without behavioral response. Thus, for a range of differences in $\mathcal{R}_{0,i}$ for the two diseases – even in the absence of interdependent behavioral feedback – the diseases may alternate in prevalence levels, as seen in Figure 3ad, an example of regime (ii). When $\mathcal{R}_{0,A}>>\mathcal{R}_{0,B}$ (Figure 3g), while the disease co-exist, disease $A$ always has higher prevalence, an example of regime (i).
Figure 3: Dynamics with three levels of spillover and three values of the basic reproduction number of disease $B$. For (a), (d) and (g), there is no spillover, $s=0$; for (b), (e), (h), there is imperfect spillover, $s=0.5$; and for (c), (f), (i), there is perfect spillover, $s=1$. For (a)-(c), $\mathcal{R}_{0,B}=2.9$; for (d)-(f), $\mathcal{R}_{0,B}=2$; and for (g)-(i) $\mathcal{R}_{0,B}=1.3$). Other parameters as in Table 1 with $\mathcal{R}_{0,A}=3$.
@@
When there is behavioral spillover, disease $A$ and disease $B$ impact each other’s spread. This alters the region of co-existence of both diseases. Under our assumption of $\mathcal{R}_{0,A}>\mathcal{R}_{0,B}>1$, disease $A$ always persists but the region where disease $B$ also persists is determined by the combination of spillover ($0<s\leq 1$) and the transmission capacity of disease $B$, as determined by $\mathcal{R}_{0,B}$. The impact of disease $A$ on disease $B$ potentially leads to its extinction even when $\mathcal{R}_{0,B}>1$ (e.g., Figure 3fi), examples of regime (iii).
We conduct a more systematic analysis, varying values of $\mathcal{R}_{0,B}$ and $s$ in small increments, and examine the final simulation results for outcome regimes. Specifically, $\mathcal{R}_{0,B}$ is iterated between 1 and 3 (with $\mathcal{R}_{0,A}=3$) and $s$ is iterated between 0 (no spillover) and 1 (perfect spillover). Figure 4 shows simulated outcomes over the first year of an epidemic for various combinations of the two parameter values. As seen in Figure 4a, as $\mathcal{R}_{0,B}$ increases, a larger amount of spillover (larger $s$) is required for exclusion of disease $B$. In Figure 4b, we quantify the percentage of time in the first year where the prevalence of disease $B$ exceeds that of disease $A$. The specified nine points marked in Figure 4 correspond to the scenarios presented in Figure 3. Considering the entire range of possible spillover and basic reproduction numbers of disease $B$ above one but below that of disease AA ($\mathcal{R}_{0,A}>\mathcal{R}_{0,B}>1$), co-existence is always possible for low spillover, but requires a large basic reproduction number for strain $B$ with high spillover (Figure 4a). Furthermore, only for high enough $\mathcal{R}_{0,B}$ and low enough spillover can strain $B$ rise in prevalence above strain $A$ at any point (Figure 4b).

### v2

@@ 1 Introduction
Most studies focus on modeling the spread of a single virus, such as influenza or coronavirus [11, 39, 25]; however, multiple pathogens often co-circulate, with the disease with the higher prevalence shifting over time [48]. For example, influenza and respiratory syncytial virus frequently overlap during seasonal epidemics, with higher prevalence observed in colder months [41]. In contrast, Lyme disease is more common during warmer seasons when outdoor activities increase [46], and rhinoviruses peak in the spring and fall [48]. This raises an important question: to what extent are the dynamics of multiple co-circulating pathogens independent of each other? This question became especially important during early months of the COVID-19 pandemic when a large fraction of public health capacity was devoted to treating COVID-19 patients [59].
@@ 1.1 Cross immunity
The phenomenon of cross-immunity and evolution of various strains and mutations have been explored through mathematical modeling, e.g. [6, 24, 35, 18]. Bhattacharyya et al. [6] examine cross-immunity among respiratory viruses, focusing specifically on RSV, HPIV, and hMPV. Using compartmental models, they analyze patterns of laboratory-confirmed infections from the western United States between 2002 and 2014, arguing that closely related paramyxoviruses can suppress each other’s transmission through immune responses triggered by prior infections. In foundational work, Gog and Grenfell [24], showed how cross-immunity among multiple pathogens can produce a range of complex, yet realistic behavior.
@@
1.2 Modeling Human behavior
There are various approaches to modeling the spread of infectious disease, ranging from mechanistic models—including deterministic and stochastic compartmental models [7, 3]—to curve-fitting approaches such as machine learning [34, 14]. A comprehensive study of models used during the COVID-19 pandemic by Rahmandad et al. [57] finds that mechanistic compartmental models outperform curve-fitting approaches, particularly in long-term projections. Importantly, their results show that incorporating behavioral feedback—where contact rates adjust in response to changes in prevalence or mortality—substantially improves projection accuracy.
To capture this effect, Rahmandad et al. [57] propose the SEIRb model, which extends the standard Susceptible-Exposed-Infectious-Recovered (SEIR) framework by including a risk-responsive behavioral feedback mechanism. This allows infectivity to vary dynamically as human behavior adapts to the evolving state of the disease. As a result, even relatively simple mechanistic models can outperform more complex alternatives in long-term forecasting. In a related study, LeJeune et al. [37] develop a similar framework in which infectivity is explicitly modeled as a function of disease prevalence, with a delay reflecting lagged behavioral response. Their results demonstrate how such formulations improve the ability of models to capture realistic epidemic waves across data from 50 U.S. states and Washington, D.C.
These studies illustrate the growing emphasis on incorporating human behavior into epidemic models. More broadly, LeJeune et al. [38] classify such approaches into exogenous and endogenous formulations. Exogenous approaches treat behavioral changes as externally imposed parameter shifts, whereas endogenous approaches model quantities such as contact rate or infectivity as functions of the disease state itself. Within endogenous frameworks, several extensions have been proposed. These include modeling heterogeneous compliance by dividing the susceptible population into behavioral subgroups [1, 53], as well as representing individuals as payoff maximizers who balance economic costs against infection risks [60, 20]. Additional approaches incorporate the spread of fear, evolving social norms, and adherence fatigue [19, 55, 52], or use game-theoretic formulations to capture strategic behavioral adaptation [27, 9].
While these works highlight recent advancements in the consideration of cross-immunity and human behavior in epidemiological models, we emphasize here a critical gap which persists in the literature. On one hand, models of multi-pathogen dynamics often lack behavioral mechanisms; on the other hand, many behavioral epidemic models focus on single-pathogen contexts, overlooking how individual and collective responses influence outbreak dynamics. Particularly, compliance with NPIs can simultaneously influence the spread of multiple pathogens with similar transmission mechanisms.
@@ 1.3 Appearance in data
In real-world settings, public behavior is rarely pathogen-specific. For instance, increasing mask-wearing or social distancing in response to one disease can inadvertently suppress the transmission of others. This was clearly observed during the COVID-19 pandemic, when widespread NPI substantially reduced flu cases [30, 62, 54]. As shown in Figure 1, the United States has historically experienced annual flu seasons, typically peaking in late fall or early winter, until the onset of the recent COVID-19 pandemic. During the surge in COVID-19 cases in 2020 and 2021, influenza cases dramatically declined, with very few cases of Influenza A or B. Interestingly, as COVID-19 was brought under control and transitioned into an endemic phase, seasonal flu patterns re-emerged. In fact, Flu A cases surged during the 2022–2023 season with a peak that was not only substantially larger but also arrived earlier than those observed in pre-pandemic years. This pattern is often attributed to behavioral factors: compliance with NPIs not only curbed COVID-19 transmission but also suppressed flu outbreaks. Once such measures were relaxed, flu incidence rebounded. This heightened rebound can be partly attributed to a decline in population immunity, as the prolonged suppression of influenza circulation during the pandemic left a larger-than-usual proportion of the population with reduced or waning protection against circulating strains [10].
Both COVID-19 and influenza have similar transmission mechanisms which can be targeted by comparable prevention measures, likely contributing to the reduction in influenza cases observed during the initial outbreak of COVID-19. However, examining the literature and data of other infectious diseases with distinct transmission routes, such as COVID-19 and Lyme disease, we see that changes in behavior in response to COVID-19 may have also caused changes in the spread of non-respiratory diseases. For example, the lifestyle disruption that COVID-19 caused across the world resulted in changes in time spent outdoors, affecting potential for human exposure to tick-borne diseases [borșan2021recreational, 29, 44]. Some regions reported an increase in tick-borne encephalitis, possibly due to more time spent out-of-doors [31]. Other regions, though, saw a reduction in reported cases of Lyme disease, potentially resulting from disruptions in the healthcare workflow and healthcare-seeking habits which led to under-reporting of cases [29, 44, 51]. In either case, these correlations further support the hypothesis that disruptions in habits and routines - i.e., changes in public behavior - in response to one disease can affect the spread of another disease with an independent transmission pathway.
@@ 1.4 Beyond specific diseases
Similar dynamics are likely to arise in a variety of contexts and among pathogens with comparable transmission characteristics. For example, when individuals increase hand washing in response to a surge in one particular disease, they may also inadvertently reduce the transmission of other pathogens that spread through similar pathways. As such, modeling frameworks that incorporate behavioral coupling between co-circulating pathogens can offer unique insights into epidemic dynamics. These models can help answer the following important questions: How do general patterns of disease spread emerge when multiple pathogens are active and NPIs targeting one also reduce the transmission of another? How might these patterns change if such spillover effects are imperfect or only partially suppress the spread of a second disease? Can we identify ranges of reproductive numbers where one pathogen exceeds the other or where multiple pathogens can coexist? Addressing these questions is essential for improving our understanding of—and response to—the complex interplay between human behavior and the concurrent spread of multiple infectious diseases.
In this paper, we study the short-term and long-term dynamics of two pathogens spreading simultaneously, where NPIs targeting one pathogen can indirectly influence the spread of others—a phenomenon we term behavioral spillover. Generally, real-world infectious disease dynamics are influenced by many time-varying factors, including seasonality, policy changes, behavioral responses, emerging variants, environmental factors, and vaccination campaigns. The objective of this study, however, is not to reproduce the full complexity of epidemic dynamics, but rather a theoretical investigation: to isolate and analyze a specific mechanism—behaviorally mediated interactions between diseases—within a simple and analytically tractable framework. To that end, we present a mathematical model of two interacting pathogens and analyze their outbreak dynamics. The interaction is solely via human behavior; that is, no cross-immunity is considered, but we model how human risk response to one can fully or partially affect the spread of the other pathogen. We postulate that there are two regimes – persistence by only one pathogen or coexistence of both – and examine tipping points for change in patterns. This equilibrium is also influenced by the basic reproduction numbers of the pathogens. Over time, we observe a likelihood of periodic shift in which disease has higher prevalence between pathogens.
@@ 2 Materials and Methods
The core of our model structure is the simple SIR model with two modifications that include behavioral response and waning immunity. We refer to this family of models as SIRSb models to indicate the return of those in the recovered compartment to the susceptible compartment, and the inclusion of behavioral feedback that affects infectivity [37].
@@
where $S_{i}$, $I_{i}$, and $R_{i}$ represent the proportion of individuals in the total population who are susceptible to disease $i$, infectious with disease $i$, and recovered from disease $i$, respectively. Notice that there is overlap among the two diseases in the population. The susceptible, infected, and recovered, populations with disease $A$ respectively, encompass the entire population, as do the susceptibles, infected, and recovered individuals with disease $B$. This stems from the assumption that all individuals have the potential to be infected with disease $A$, disease $B$, or both diseases simultaneously; in other words, individuals can become infected with both diseases without any biological cross-immunity. For example, a disease $A$ infectious individual is also simultaneously one of susceptible, infectious or recovered from disease $B$. Furthermore, infectious individuals do not change their behavior upon infection (e.g., no quarantine or isolation is included).
Susceptible individuals become infected with disease $i$ through interactions with infectious individuals of disease $i$ with transmission rate $\beta_{i}(\cdot)$, which can depend on other variables. Infectious individuals recover after an average period of $\tau_{I}$ days. Once recovered, individuals lose immunity and are once again susceptible after on average $\tau_{R}$ days.
Here, $\widetilde{I}_{i}$ represents the perceived number of disease $i$ infectious individuals, which follows the true value $I_{i}$ with an average delay $\tau_{P}$. Specifically, $\widetilde{I}_{i}$ adjusts toward $I_{i}$ at rate $1/\tau_{P}$, so that discrepancies between perceived and actual infections decay over time [47, 64]. This formulation corresponds to a standard exponential smoothing (information delay) process previously used in the literature to represent information flow (e.g., [65, 40]).
@@
which represents the effect of response (e.g., decline in contact rate) on the transmission rate of $\beta_{i}$. Here, $m_{i}$ is formulated such that increases in prevalence ($I$), tracked through a lagged compartment of the infectious population, $\widetilde{I}_{i}$, lead to a decline in $m_{i}$. As prevalence $I_{i}$ increases, perceived risk of $\widetilde{I}_{i}$ (modeled by an exponential delay of prevalence) also increases. Higher perceived risk leads to stronger reaction, decreasing $m_{i}$, and consequently decreasing $\beta_{i}$.
@@ 2.1 Transmission rate
We consider three scenarios that differ with respect to the level of behavioral spillover. Mathematically, this leads to differences in the multipliers $m_{A}$ and $m_{B}$ in the formulation of the transmission rate, denoted by $\beta_{i}(\cdot)$, in the SIRS model given in System (LABEL:eqn:SIRS). The specific system of equations for each scenario, with the relevant formulation of transmission, is included in Appendix A1.1 for the reader’s reference.
@@ 2.1.1 Two independent diseases
For the case of two pathogens that are not behaviorally coupled and with no cross-immunity, we assume that the transmission rate of each pathogen is formulated independently as
@@
where $\beta_{0,i}$ is the infectivity of disease $i$ and $m_{i}$ the multiplier of disease $i$. In this case, disease $A$ and disease $B$ are spreading independently.
@@ 2.1.2 Two behaviorally-coupled diseases
As a first scenario, we assume the simple case of perfect spillover. This represents a situation where a proportional decline in transmission of disease $A$, due to behavioral response, results in the same proportional decline in the transmission of disease $B$, and vice versa. This happens when the two diseases transmit in an identical manner. Behavioral response to one can alter the spread of the other, such that transmission of each disease is affected by both $m_{A}$ and $m_{B}$. Thus, we modify the transmission rate of disease $i$, $\beta_{i}$, to be
@@
It is important to note that this formulation encapsulates both no spillover (independent diseases) and perfect spillover. For perfect spillover ($s=1$), the overall effect of the response to $B$ on the spread of disease $A$ is $m_{B}$, and for no spillover ($s=0$), the overall effect of response to disease $B$ on the spread of $A$ is constant at one. Table 1 summarizes all the formulations for $\beta_{i}(\cdot)$ with $i\in\{A,B\}$. We note that in model formulation and analytical results, we present no spillover ($s=0$) and perfect spillover ($s=1$) before discussing imperfect spillover ($0<s<1$), while for numerical results, we present results ordered according to increasing spillover.
@@ 3 Results
We begin by numerically simulating the dynamics of our system with a range of spillover – none, perfect and imperfect (Section 3.1). We choose parameters from the literature on COVID-19 and influenza to highlight our application of this model to concrete scenarios. The parameter values used for analysis and simulation of the two diseases ($A$ and $B$) are given in Table 2. Following our numerical examination, we determine the existence and stability of equilibria of the model under each spillover scenario (Section 3.2) and assess the structural and practical identifiability of the model (Section 3.3).
@@ 3.1 Numerical results
To examine the dynamics of our behaviorally-coupled system, we numerically simulate our system in MATLAB R2021a [28] using ode45 and observe the dynamics over short (one year) and long (multiple years) time scales. Figure 3 shows results from nine simulation experiments which vary the level of spillover and the basic reproduction number of disease $B$ (with the basic reproduction number of disease $A$ fixed at 3). A helpful quantity for explaining the behavior in the system is the effective reproduction number, $\mathcal{R}_{e}$, which tells if the disease is expanding or contracting at the current time. For our two disease system, we define the effective reproduction number of disease $i$, $\mathcal{R}_{e,i}$, as
@@
Symbol Description Value Source
$\mathcal{R}_{0,A}$ basic reproduction number of disease AA 3 [13]
$\mathcal{R}_{0,B}$ basic reproduction number of disease $B$ [1,3] [49]
$\beta_{0,A}$ infectivity rate of disease AA $\frac{\mathcal{R}_{0,A}}{\tau_{I}}$ [13, 8]
$\beta_{0,B}$ infectivity rate of disease $B$ $\frac{\mathcal{R}_{0,B}}{\tau_{I}}$ [49, 8]
$\tau_{I}$ infection period 7 days [8]
$\tau_{P}$ delay to adjust risk perception 30 days [57]
$\tau_{R}$ immunity period 100 days [63]
$k$ sensitivity to risk 100 [23]
$s$ spillover constant 0 varied in [0,1]
$S_{i}(0)$ initial proportion of disease-$i$ susceptible individuals 0.9999 assumed
$I_{i}(0)$ initial proportion of disease-$i$ infectious individuals 0.0001 assumed
$R_{i}(0)$ initial proportion of disease-$i$ recovered individuals 0 assumed
$\widetilde{I}_{i}(0)$ initial proportion of disease-$i$ perceived infectious individuals 0 assumed
Without spillover (under our assumption $\mathcal{R}_{0,A}>\mathcal{R}_{0,B}>1$), there are two regimes of behavior: (i) diseases $A$ and $B$ co-exist but disease $A$ prevalence always remains above disease $B$ prevalence and (ii) diseases $A$ and $B$ co-exist but the diseases alternate whose prevalence is higher. As depicted in the left column of Figure 3 (Panels 3(a), 3(d), 3(g)), both diseases persist as the diseases are spreading independently and (by assumption) the basic reproduction number of each disease is above one. As the diseases are not interacting in any way, the endemic level of the disease is dictated solely by its basic reproduction number. Furthermore, the dynamics are qualitatively similar for the two diseases and are quantitatively identical when the diseases have the same traits. As cases grow for disease $i$, individuals respond, leading to reductions in the transmission rate and a drop in $\mathcal{R}_{e,i}$. Ultimately, $\mathcal{R}_{e,i}$ settles to one such that cases reach a non-zero equilibrium (under our assumption $\mathcal{R}_{0,i}>1$). As the basic reproduction number of disease $B$ is lowered, the initial peak occurs later and reaches a lower level (Figure 3adg, red lines as $\mathcal{R}_{0,B}$ is lowered from 2.9 to 2 to 1.3). This is because the natural disease infectivity is lower, reducing the transmission rate even without behavioral response. Thus, for a range of differences in $\mathcal{R}_{0,i}$ for the two diseases – even in the absence of interdependent behavioral feedback (i.e., diseases spread entirely independently) – the diseases may alternate in prevalence levels, as seen in Figure 3ad, an example of regime (ii). When $\mathcal{R}_{0,A}>>\mathcal{R}_{0,B}$ (Figure 3g), while the disease co-exist, disease $A$ always has higher prevalence, an example of regime (i).
Figure 3: Dynamics with three levels of spillover and three values of the basic reproduction number of disease $B$. For (a), (d) and (g), there is no spillover, $s=0$, which corresponds to the two diseases spreading independently (i.e., the curve is identical to that of the single disease spreading in the population); for (b), (e), (h), there is imperfect spillover, $s=0.5$; and for (c), (f), (i), there is perfect spillover, $s=1$. For (a)-(c), $\mathcal{R}_{0,B}=2.9$; for (d)-(f), $\mathcal{R}_{0,B}=2$; and for (g)-(i) $\mathcal{R}_{0,B}=1.3$). Other parameters as in Table 2 with $\mathcal{R}_{0,A}=3$.
@@
When there is behavioral spillover, disease $A$ and disease $B$ impact each other’s spread. This alters the region of co-existence of both diseases. Under our assumption of $\mathcal{R}_{0,A}>\mathcal{R}_{0,B}>1$, disease $A$ always persists but the region where disease $B$ also persists is determined by the combination of spillover ($0<s\leq 1$) and the transmission capacity of disease $B$, as determined by $\mathcal{R}_{0,B}$. The impact of disease $A$ on disease $B$ potentially leads to its extinction even when $\mathcal{R}_{0,B}>1$. For example, regime (iii) is seen in Figure 3fhi.
We conduct a more systematic analysis, varying values of $\mathcal{R}_{0,B}$ and $s$ in small increments, and examine the final simulation results for outcome regimes. Specifically, $\mathcal{R}_{0,B}$ is iterated between 1 and 3 (with $\mathcal{R}_{0,A}=3$) and $s$ is iterated between 0 (no spillover) and 1 (perfect spillover). Figure 4 shows simulated outcomes over the first year of an epidemic for various combinations of the two parameter values. As seen in Figure 4a, as $\mathcal{R}_{0,B}$ increases, a larger amount of spillover (larger $s$) is required for exclusion of disease $B$. In Figure 4b, we quantify the percentage of time in the first year where the prevalence of disease $B$ exceeds that of disease $A$. However, when disease $B$ prevalence exceeds disease $A$, it may not be by much. Thus, in Figures 4c and d we quantify the cumulative amount across one year by which one disease exceeds the other. The specified nine points marked in Figure 4 correspond to the scenarios presented in Figure 3. Considering the entire range of possible spillover and basic reproduction numbers of disease $B$ above one but below that of disease AA ($\mathcal{R}_{0,A}>\mathcal{R}_{0,B}>1$), co-existence is always possible for low spillover, but requires a large basic reproduction number for strain $B$ with high spillover (Figure 4a). Furthermore, only for high enough $\mathcal{R}_{0,B}$ and low enough spillover can strain $B$ rise in prevalence above strain $A$ at any point (Figure 4b) and the level by which it exceeds disease $A$ may be quite low (Figure 4d). Our results on the impact of behavioral spillover are robust to another choice of the behavioral feedback, as demonstrated by the use of a fractional rather than exponential formulation (see [38] for discussion of various functional formulations for behavioral feedback). In particular, we reproduce Figs. 3 and 4 in Section A2 with a discussion of the fractional formulation.